OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that passes through (2,5) and is parallel to the graph of y=x-3. Help! help! I keep getting 2 for b but the answer book says its 3. walk me through this!!!

OpenStudy (vampirediaries):

Do you know the equation for the slope-intercept form?

OpenStudy (anonymous):

Once you do know the slope-intercept form, note that another line parallel to it will have the same slope but a different intercept.

OpenStudy (j2lie):

y-5=1(x-5) y=x+5. The y-intercept have to be different and it doesn't have have the same sign.